OpenStudy (anonymous):

Solve the initial value first order linear differential equation: y'=2y+x & y(0)=1

OpenStudy (anonymous):

@TuringTest

OpenStudy (anonymous):

@mahmit2012

OpenStudy (valpey):

\[y'-2y=x\] \[y_c'-2y_c=0;\ and\ y_p'-2y_p=x\] \[m=2\ so\ y_c=Ae^{2x}\] \[Let\ y_p=C_1x+C_2\] \[Then\ y_p'=C_1\] \[And\ C_1-2(C_1x+C_2)=x\] \[So\ C_1-2C_2=0\ \ and \ -2C_1=1\]

OpenStudy (valpey):

Follow me? \[y=y_c+y_p\]

OpenStudy (anonymous):

Wouldn't it be Ae^-2x? Also, how do you find m?

OpenStudy (valpey):

No. If it were Ae^(-2x) then \[-2Ae^{-2x}-2Ae^{-2x}=0\] As for m: \[y_c'−2y_c=0\] \[m=\frac{y_c'}{y_c}\] \[m-2=0\]

OpenStudy (anonymous):

So would it be: \[y=-x/2-1/4+(3/4)e ^{-2x}\]

OpenStudy (valpey):

Almost. I think A would need to be 5/4 for y(0) to be 1.

OpenStudy (anonymous):

Thanks

OpenStudy (valpey):

And again \[Ae^{2x}\ not\ Ae^{-2x}\]

OpenStudy (anonymous):

Got it

OpenStudy (valpey):

Sweet.